# Doomed Worlds I: No new evidence for orbital decay in a long-term survey of 43 ultra-hot Jupiters

### v1

@@ 7 Preliminary population analysis: where are the decaying Ultra-hot Jupiters?
7.1 Observational constraints: in what systems could we detect even WASP-12 b like decay?
@@
Should we expect to have detected orbital decay in a substantial number of systems? To answer this question, we performed two simple tests to determine if we could have even detected decay as rapid as that of WASP-12 b. Theoretically, we would expect a wide range of decay rates commensurate with the diverse stellar and planetary properties for these systems (see the range of predicted values in Table 7), but using WASP-12 b’s relatively rapid rate sets a useful baseline against which to test our ability to detect tidal decay. The results of these tests are listed in Table 8. The first test assigns all planets the same rate of decay as WASP-12 b, ms yr (Yee et al., 2020). We find that nearly half (18 out of 42 planets that are not WASP-12 b) would have had , as shown in the Test 1 column in Table 8, while in reality only one did – and that planet, WASP-121 b, has a positive value for . Four systems (WASP-19 b, WASP-43 b, WASP-18 b and TrES-3 b) would have values up to ten times larger than that of WASP-12 b if they were decaying at the same rate, making the continued non-detection of decay in these systems highly significant.
@@
Thus, for 20 out of 42 systems, there is evidence that we could have detected WASP-12 b like decay, but have not, meaning these planets are likely to not be decaying as rapidly. The flip side of this statement is that for the other half of our sample (22/42 planets), we still do not have enough baseline to detect even decay as rapid as WASP-12 b. With the shortest of these planets having baselines that extend just 4 years, it will take several more years to perhaps even decades to determine just how unusual WASP-12 b is.
@@ 7.3 Comparison to theory
The resulting theoretical estimates for are listed as in Table 3, and we computed the theoretical decay time and as in Eq. 15 in Weinberg et al. (2024). For stellar models with radiative cores and convective envelopes, Weinberg et al. (2024) found values that range from (at ) to at . For higher-mass stars () on the main sequence, the cores are convective and in such cases is likely to be very large; for all such stars, we set as a representative value. If those stars were instead subgiants, however, as might be the case for WASP-12 b, their would be much smaller since their cores would become radiative.
@@
| TOI-2109b | -4.5 | – | ||
| a. Test 1: analytical BIC if WASP-12 b like decay ( ms y) | ||||
@@ 8.1 No evidence yet for orbital decay being a common UHJ trait
We note that of the two systems for which orbital decay is a credible explanation, both Kepler-1658 b (Vissapragada et al., 2022) and WASP-12 b (Weinberg et al., 2017) may have evolved off the main sequence, which could offer an explanation for more rapid decay (e.g., Weinberg et al., 2024). However, the status of the star WASP-12 is unknown, with a main sequence star still favored (Bailey & Goodman, 2019). If WASP-12 is in fact evolved, that would leave no main sequence stars known to host decaying planets, offering support to predictions that orbital decay is unlikely during the main sequence and is a feature of later stages of stellar life. This is however in tension with the independent observation that hot Jupiter host stars are younger than all planet hosts (Hamer & Schlaufman, 2019; Miyazaki & Masuda, 2023; Chen et al., 2023) and that tidal decay does occur on the main sequence (Hamer & Schlaufman, 2019). It is thus important to search more UHJs both on and off the main sequence for doomed worlds, to see if this suggestive connection holds beyond a sample size of 2 planets.
@@ 8.2 Best practices for long-term timing analyses
is a powerful tool for identifying systems of interest, especially accompanied by a significant value for . However, it is only as good as the error bars of the accompanying data. With the exception of a uniform dataset for which the errors are well known (e.g., stacking all of the transits in a single Kepler quarter or TESS sector, as was done for Kepler-1658b in Figure 9), it is likely that some errors are underestimated, increasing the apparent values for . Thus, when using with real data, it is important to understand that light curves in the published literature may have mistakes in both the absolute timing and in the size of the nominal error bars. We recommend care in interpreting results, and suggest tests described in this work, including as the omit-one test to identify points that may be having an out-sized impact on the apparent result, and the rescaled test, to identify if overall errors on the system may be too small. Given our experience in this work values below 30-50 should be treated as tentative detections in need of confirmation.
@@
In the ideal case, someone would observe a transit of every planet that might be experiencing orbital decay at least once a year to avoid long gaps in the timing data, following the recommendation of Jackson et al. (2023). We note for instance that WASP-121 b, which has very few distinct epochs available, would be a much clearer case if it had been more regularly observed in the years following its first announcement. Even relatively modest aperture telescopes, such as the 14-inch used in this work, are useful since many UHJs are found around sufficiently bright stars to constrain the transit timing. Widespread networks of small scopes may also provide robust facilities for keeping ephemerides fresh and improving prospects for tidal-decay detection (Peluso et al., 2023). In fact, phasing observations of a single transit conducted by multiple, small telescopes simultaneously could result in a very high effective timing resolution, with the potential to significantly reduce mid-transit time uncertainties.
@@
There are timing errors in the literature, and the current system tends to propagate rather than eliminate errors. We identified four separate errors spanning multiple published works in the timing of CoRoT-2 b and WASP-19 b. These errors had a strong effect, in that correcting them removed the signal of tidal decay entirely; but other errors may well be lurking which conform more to our expectations, and thus have not been brought to light. We discuss this last item in detail in the next section.
@@ 8.3 The timing community has a database problem
It is becoming clear that the scientific record is currently not well set up to facilitate long-term timing searches for slow changes over time involving multiple instruments and observers over many years. Timing errors, particular conversion between times in UTC (what most data is recorded in) and BJD/TDB (what we need to use to detect true dynamical changes), are easy to make, hard to detect, and even harder to correct for posterity. Additionally, far too many early transit light curves have never been published except as figures in papers, removing valuable years of observational baseline from reanalysis. In some cases, not even a fitted midtime is available for individual light curves. And the current model of publication, whereby each new analysis publishes a list that combines midtimes of newly observed transits and times taken from the literature, has led to compilations of compilations of midtimes. We have documented multiple cases where this process has introduced errors in one work that are then propagated through multiple works.
@@
The problem of data curation currently constitutes a major component of all long-term timing analyses. Rather than each group repeating the same effort to identify, extract, convert, reformat, and refit prior transit light curve data, there should be a single archival database that contains a definitive list of all prior published transit times, as taken directly from their original publication. If more than one work has reported a time for a given transit, it should be clear which times are re-analyses, and a best time should be indicated, similar to how the NASA Exoplanet Archive reports all values for stellar and planetary parameters but also provides a best value for researchers to use. This ideal transit-timing database would provide information about each individual light curve, and make transit light curves available for download in a uniform format to make it easy to re-fit data that is supposed to be in the public domain, but in practice has been inconvenient to access. The Exoplanet Transit Database is an excellent resource and is well used by amateurs, but it is missing some key features (notably, it uses HJD/UTC instead of BJD/TDB timing and lacks a bulk download option for both light curves and the original JD times before conversion) and moreover has low uptake from the professional community. It would require either a substantial overhaul of the ETD or an entirely new system to be developed and funded for the long term.
@@
In addition to observations of primary transits, an ideal timing database would also include timing constraints placed by planetary occultations (also called secondary eclipses). Occultations are even less likely to have published timing data available for individual light curves, owing to both their lower signal to noise and to the propensity of occultation data to be used to study planetary atmospheres, where stacked and phased data of multiple occultations are necessary to tease out small signals. Occultation timing observations are used to determine if apsidal precession could explain a timing trend, since orbital decay will cause the interval between both transits and occultations to decrease, while precessing systems would have opposite trends, with occultations showing increasing periods while transits show decreasing periods, and vice versa.
@@ Acknowledgments
ERA and AAS were visiting Astronomers at the Infrared Telescope Facility, which is operated by the University of Hawaii under contract 80HQTR19D0030 with the National Aeronautics and Space Administration. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain, and we recognize that we are guests.
@@
We respectfully acknowledge the traditional custodians of all lands throughout Australia, and recognise their continued cultural and spiritual connection to the land, waterways, cosmos, and community. We pay our deepest respects to all Elders, ancestors and descendants of the Giabal, Jarowair, and Kambuwal nations, upon whose lands the Minerva-Australis facility at Mt Kent is situated.

### v2

@@ 7 Preliminary Population Analysis: Where Are the Decaying UHJs?
7.1 Observational Constraints: In What Systems Could We Even Detect WASP-12 b-Like Decay?
@@
Should we expect to have detected orbital decay in a substantial number of systems? To answer this question, we performed two simple tests to determine if we could have even detected decay as rapid as that of WASP-12 b. Theoretically, we would expect a wide range of decay rates commensurate with the diverse stellar and planetary properties for these systems (see the range of predicted values in Table 7), but using WASP-12 b’s relatively rapid rate sets a useful baseline against which to test our ability to detect tidal decay. The results of these tests are listed in Table 8. The first test assigns all planets the same rate of decay as WASP-12 b, ms yr -1 (Yee et al., 2020). We find that nearly half (18 out of 42 planets that are not WASP-12 b) would have had , as shown in the Test 1 column in Table 8, while in reality only one did – and that planet, WASP-121 b, has a positive value for . Four systems (WASP-19 b, WASP-43 b, WASP-18 b and TrES-3 b) would have values up to ten times larger than that of WASP-12 b if they were decaying at the same rate, making the continued non-detection of decay in these systems highly significant.
@@
Thus, for 20 out of 42 systems, there is evidence that we could have detected WASP-12 b like decay, but have not, meaning these planets are likely to not be decaying as rapidly. The flip side of this statement is that for the other half of our sample (22/42 planets), we still do not have enough baseline to even detect decay as rapid as WASP-12 b. With the shortest of these planets having baselines that extend just 4 years, it will take several more years to perhaps even decades to determine just how unusual WASP-12 b is.
@@ 7.3 Comparison to theory
The resulting theoretical estimates for are listed as in Table 3, and we computed the theoretical decay time and rate of period decrease as in Eq. 15 in Weinberg et al. (2024). For stellar models with radiative cores and convective envelopes, Weinberg et al. (2024) found values that range from (at ) to at . For higher-mass stars () on the main sequence, the cores are convective and in such cases is likely to be very large; for all such stars, we set as a representative value. If those stars were instead subgiants, however, as might be the case for WASP-12 b, their would be much smaller since their cores would become radiative.
@@
| TOI-2109b | -4.5 | – | c | |
| a. Test 1: analytical BIC if WASP-12 b like decay ( ms y-1) (Yee et al., 2020) | ||||
@@ 8.1 No Evidence Yet for Orbital Decay Being a Common UHJ Trait
We note that of the two systems for which orbital decay is a credible explanation, both Kepler-1658 b (Vissapragada et al., 2022) and WASP-12 b (Weinberg et al., 2017) may have evolved off the main sequence, which could offer an explanation for more rapid decay (e.g., Weinberg et al., 2024). However, the status of the star WASP-12 is ambiguous (Leonardi et al., 2024), with a main sequence star still favored (Bailey & Goodman, 2019). If WASP-12 is in fact evolved, that would leave no main sequence stars known to host decaying planets, offering support to predictions that orbital decay is unlikely during the main sequence and is a feature of later stages of stellar life. This is however in tension with the independent observation that hot Jupiter host stars are younger than all planet hosts (Hamer & Schlaufman, 2019; Chen et al., 2023; Miyazaki & Masuda, 2023) and that tidal decay does occur on the main sequence (Hamer & Schlaufman, 2019). It is thus important to search more UHJs both on and off the main sequence for doomed worlds to see if this suggestive connection holds beyond a sample size of two planets.
@@ 8.2 Best Practices for Long-term Timing Analyses
is a powerful tool for identifying systems of interest, especially accompanied by a significant value for . However, it is only as good as the error bars of the accompanying data. With the exception of a uniform dataset for which the errors are well known (e.g., stacking all of the transits in a single Kepler quarter or TESS sector, as was done for Kepler-1658b in Figure 10), it is likely that some errors are underestimated, increasing the apparent values for . Thus, when using with real data, it is important to understand that light curves in the published literature may have mistakes in both the absolute timing and in the size of the nominal error bars. We recommend care in interpreting results and suggest the tests described in this work, including the omit-one test to identify points that may be having an out-sized impact on the apparent result, and the rescaled test to identify if overall errors on the system may be too small. Given our experience in this work values below 30-50 should be treated as tentative detections in need of confirmation.
@@
In the ideal case, someone would observe a transit of every planet that might be experiencing orbital decay at least once a year to avoid long gaps in the timing data, following the recommendation of Jackson et al. (2023). We note for instance that WASP-121 b, which has very few distinct epochs available, would be a much clearer case if it had been more regularly observed in the years following its first announcement. Even relatively modest aperture telescopes, such as the 14-inch used in this work, are useful since many UHJs are found around sufficiently bright stars to constrain the transit timing. Widespread networks of small scopes may also provide robust facilities for keeping ephemerides fresh and improving prospects for tidal-decay detection (Peluso et al., 2023). In fact, phasing observations of a single transit conducted by multiple small telescopes simultaneously could result in a very high effective timing resolution, with the potential to significantly reduce mid-transit time uncertainties.
@@
There are timing errors in the literature, and the current system tends to propagate rather than eliminate errors. We identified errors spanning multiple published works in the timing of CoRoT-2 b and WASP-19 b. These errors had a strong effect, in that correcting them removed the signal of tidal decay entirely; but other errors may well be lurking that conform more to our expectations and thus have not been brought to light. We discuss this last item in detail in the next section.
@@ 8.3 The Timing Community Has a Database Problem
It is becoming clear that the scientific record is currently not well set up to facilitate long-term timing searches for slow changes over time involving multiple instruments and observers over many years. Timing errors, particular conversion between times in UTC (what most data are recorded in) and BJD/TDB (what we need to use to detect true dynamical changes), are easy to make, hard to detect, and even harder to correct for posterity. Additionally, far too many early transit light curves have never been published except as figures in papers, removing valuable years of observational baseline from reanalysis. In some cases, not even a fitted midtime is available for individual light curves. And the current model of publication, whereby each new analysis publishes a list that combines midtimes of newly observed transits and times taken from the literature, has led to compilations of compilations of midtimes. We have documented multiple cases where this process has introduced errors in one work that are then propagated through multiple works.
@@
The problem of data curation currently constitutes a major component of all long-term timing analyses. Rather than each group repeating the same effort to identify, extract, convert, reformat, and refit prior transit light-curve data, there should be a single archival database that contains a definitive list of all prior published transit times, as taken directly from their original publication. If more than one work has reported a time for a given transit, it should be clear which times are re-analyses, and a best time should be indicated, similar to how the NASA Exoplanet Archive reports all values for stellar and planetary parameters but also provides a best value for researchers to use. This ideal transit-timing database would provide information about each individual light curve, and make transit light curves available for download in a uniform format to make it easy to re-fit data that is supposed to be in the public domain, but in practice has been inconvenient to access. The Exoplanet Transit Database is an excellent resource and is well used by amateurs, but it is missing some key features (notably, it uses HJD/UTC instead of BJD/TDB timing and lacks a bulk download option for both light curves and the original JD times before conversion) and, moreover, has low uptake from the professional community. It would require either a substantial overhaul of the ETD or an entirely new system to be developed and funded for the long term.
@@
In addition to observations of primary transits, an ideal timing database would also include timing constraints placed by planetary occultations (also called secondary eclipses). Occultations are even less likely to have published timing data available for individual light curves, owing to both their lower signal-to-noise and to the propensity of occultation data to be used to study planetary atmospheres, where stacked and phased data of multiple occultations are necessary to tease out small signals. Occultation timing observations are used to determine if apsidal precession could explain a timing trend, since orbital decay will cause the interval between both transits and occultations to decrease, while precessing systems would have opposite trends, with occultations showing increasing periods while transits show decreasing periods, and vice versa.
@@ Acknowledgments
E.R.A. and A.A.S. were visiting astronomers at the Infrared Telescope Facility, which is operated by the University of Hawaii under contract 80HQTR19D0030 with the National Aeronautics and Space Administration. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain, and we recognize that we are guests.
@@
We respectfully acknowledge the traditional custodians of all lands throughout Australia and recognise their continued cultural and spiritual connection to the land, waterways, cosmos, and community. We pay our deepest respects to all Elders, ancestors and descendants of the Giabal, Jarowair, and Kambuwal nations, upon whose lands the Minerva-Australis facility at Mt. Kent is situated.